OpenStudy (anonymous):

Solve for the variable r in this equation: s = 2πrh i'm stumped on this and this is the last time I'm able to take this test. help is greatly appreciated thank you!!

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

There are no numbers. So best we can do is rearrange the equation. So divide both sides of equation by\[2\pi h\] So you get \[s/(2\pi h) = r\]